i do not understant how to answer this question could i get help please. i need to determine how many years it will take for the two accounts to reach the same balance. 1000 investment with a rate of 5.25% compounded continuously and 800 investment with a rate of 6.25% compounded continously. i got this far
1000e^.0525=800e^.0625

Answer by Alan3354(69443)

how many years it will take for the two accounts to reach the same balance. 1000 investment with a rate of 5.25% compounded continuously and 800 investment with a rate of 6.25% compounded continously. i got this far
1000e^.0525=800e^.0625
----------------------
1000e^.0525t=800e^.0625t You need the time in the exponents
--------------
Take the ln of both sides
ln(1000) + 0.0525t = ln(800) + 0.0625t
0.0625t - 0.0525t = ln(1000) - ln(800)
0.01t =~ 0.223144
t =~ 22.3144 years (assuming the int rates are annual)